OpenStudy (anonymous):

write a system of linear equations that relates the total cost for each printer to the number of cartridges needed. the first printer is $50 and $30 for each cartridge and the second printer is $80 and $27 for each cartridge. what is the solution to the system of equations?

OpenStudy (anonymous):

do you know what a linear equation is?

OpenStudy (anonymous):

sort of but i dont know how to write one for this problem

OpenStudy (anonymous):

ok is this a test or a homework question

OpenStudy (anonymous):

i guess it would be considered homework

OpenStudy (anonymous):

ok are there options to choose from?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

linear equations: x=50+30* the y amount of cartridges linear equations: y=80+27* the x amount of cartridges

OpenStudy (anonymous):

okay and i need to know the solutions to them too, can you help me with that part too?

OpenStudy (anonymous):

is there a total you need to find?

OpenStudy (anonymous):

i just need to know what x and y is i think

OpenStudy (anonymous):

50+30x=80+27x now what we want to do is cancel out all of the numbers that have a variable next to them

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

@breannagraber